OpenStudy (anonymous):

PLEASE HELP ME AND I WILL GIVE MEDAL AND BE SSOOOO GREATFUL <3 3.Carmen is trying to determine if the following statement is always, sometimes, or never true: The square of a prime number is odd. She writes the following statements: 32 = 9, 52 = 25, 72 = 49, 112 = 121, and 132 = 169. Using her results, she concludes that the statement is always true. a. What kind of reasoning did Carmen use? b. Is her conclusion correct? If not, use a counterexample to prove why it is not. Answer:

OpenStudy (anonymous):

@Haseeb96 @ganeshie8 @HELP!!!! @Here_to_Help15

OpenStudy (anonymous):

Please help anyone explain to me what kind of reasoning she used.

OpenStudy (anonymous):

@BassCatcher15 @study100 @cram

OpenStudy (haseeb96):

For A , it is never be true because here he said that the square of prime number is odd then in all these examples like first one 32=9 here 9 is prime number but it is equal to an even number not odd so it will never be true

OpenStudy (here_to_help15):

sorry i was watching a movie lol

OpenStudy (anonymous):

its ok lol

OpenStudy (anonymous):

ohh okay but someone told me it was true.

OpenStudy (haseeb96):

it will be always true when the square of prime number is even. for B, this conclusion is not correct according to the given statements the square of the prime number is odd but the given statement didnt satisfy it so when 31=9 ,51=25,71=49 ... these above statements satisfy the given statement( the square of the prime number is odd )

OpenStudy (haseeb96):

i have prove it

OpenStudy (anonymous):

wait did you mean 3^2=9?

jimthompson5910 (jim_thompson5910):

2 is a prime number but 2 squared = 2*2 = 4 which is even (and not odd)

jimthompson5910 (jim_thompson5910):

so this counter-example shows you that her claim is sometimes true and not always true

OpenStudy (haseeb96):

9 is the prime number and 32 is an even it said that when u square the prime number then odd will come 31=9

OpenStudy (anonymous):

oh okay. So wait what do I put for A her reasoning?@jim_thompson5910

jimthompson5910 (jim_thompson5910):

well she's just listing a few examples. She hasn't proved it for ALL prime numbers. Just for this small list.

jimthompson5910 (jim_thompson5910):

So her reasoning is flawed in that she hasn't accounted for every prime

OpenStudy (anonymous):

Oh okay thank you! :)

jimthompson5910 (jim_thompson5910):

np